OpenStudy (anonymous):

What is the projection of (3,4) onto (3,1)? (Points : 2) 1.3(3,1) 0.7(3,1) 1.3(2,4) 0.7(2,4) according to the lesson it would be something like : vw/w but i think i would still be using a unit circle can someone help with this???

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

are they two vectors?

OpenStudy (anonymous):

yes, they are so far i have got (9,4)/(3,1 = 9 + 4 / sqrt4

OpenStudy (michele_laino):

the requested projection, is given by the subsequent formula: if A=(3,4), B=(3,1), then, we have: projection= \[\large \frac{{A \cdot B}}{{B \cdot B}}B\]

OpenStudy (michele_laino):

where I used the canonical scalar product

OpenStudy (michele_laino):

so, substituting your data, I get: \[\large \frac{{A \cdot B}}{{B \cdot B}}B = \frac{{9 + 4}}{{9 + 1}}\left( {3,1} \right) = ...?\]

OpenStudy (anonymous):

sorry had to open the door for my dad you would get : 13/10 (3,1)

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

and 13/10 = 1.3

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

i guess i was on the wrong page in the lesson then lol the aswer would be 1.3 (3,1)

OpenStudy (anonymous):

just to be sure i read this right in the lesson: If the dot product of two nonzero vectors v1 and v2 is zero, what does this tell us? (Points : 2) v1 is perpendicular to v2. v1 is a component of v2. v1 is parallel to v2. v1 = v2 the answer would be a right because i have c selected and i dont know why...

OpenStudy (michele_laino):

by definition is the first option, namely v_1 and v_2 are perpendicular each other

OpenStudy (anonymous):

ok, lol thought that wasn't right, thank you for the help! :)

OpenStudy (michele_laino):

thank you! :)